OpenStudy (anonymous):

Prove that: cos alpha/a +cos beta / b +cos gamma / c = a^2+b^2+c^2/2abc

OpenStudy (anonymous):

\[\frac{ \cos \alpha }{ a }+\frac{ \cos \beta }{ b }+\frac{ \cos \gamma }{ c }=\frac{ a^2+b^2+c^2 }{ 2abc }\]

OpenStudy (anonymous):

Reference:- \[\cos \alpha=\frac{ b^2+c^2-a^2 }{ 2bc }\]

OpenStudy (anonymous):

\[\cos \beta=\frac{ a^2+c^2-b^2 }{ 2ac }\]

OpenStudy (anonymous):

\[\cos \gamma=\frac{ a^2+b^2-c^2 }{ 2ab }\]

OpenStudy (anonymous):

Using L.H.S :- \[\frac{ \cos \alpha }{ a }+\frac{ \cos \beta }{ b }+\frac{ \cos \gamma }{ c }\]

OpenStudy (raden):

so, (b^2+c^2-a^2)/(2abc)+(a^2+c^2-b^2)/(2abc)+(a^2+b^2-c^2)/(2abc) (a^2+b^2+c^2)/(2abc) proof ^_^

OpenStudy (anonymous):

0_0

OpenStudy (anonymous):

whoa!! that was easy!!

OpenStudy (raden):

yes, its very easy :)